# EgoPoseFormer: A Simple Baseline for Egocentric 3D Human Pose Estimation

### v1

@@ Abstract
We present EgoPoseFormer, a simple yet effective transformer-based model for stereo egocentric human pose estimation. The main challenge in egocentric pose estimation is overcoming joint invisibility, which is caused by self-occlusion or a limited field of view (FOV) of head-mounted cameras. Our approach overcomes this challenge by incorporating a two-stage pose estimation paradigm: in the first stage, our model leverages the global information to estimate each joint’s coarse location, then in the second stage, it employs a DETR style transformer to refine the coarse locations by exploiting fine-grained stereo visual features. In addition, we present a Deformable Stereo Attention operation to enable our transformer to effectively process multi-view features, which enables it to accurately localize each joint in the 3D world. We evaluate our method on the stereo UnrealEgo dataset and show it significantly outperforms previous approaches while being computationally efficient: it improves MPJPE by 27.4mm (45% improvement) with only 7.9% model parameters and 13.1% FLOPs compared to the state-of-the-art. Surprisingly, with proper training techniques, we find that even our first-stage pose proposal network can achieve superior performance compared to previous arts. We also show that our method can be seamlessly extended to monocular settings, which achieves state-of-the-art performance on the SceneEgo dataset, improving MPJPE by 25.5mm (21% improvement) compared to the best existing method with only 60.7% model parameters and 36.4% FLOPs.
@@ 1 Introduction
In the rapid expansion of Virtual Reality (VR) and Augmented Reality (AR) technologies [1, 2], the capability to accurately interpret and emulate human actions becomes increasingly crucial. Central to this pursuit is the egocentric pose estimation task [11, 31, 12, 37, 36, 3, 32, 35, 26, 41, 22, 45, 14], which aims to estimate the 3D body pose from a vantage point inherent to the user, predominantly from head-mounted cameras. Its precision is pivotal in VR/AR applications, e.g., gaming and virtual meeting, rendering it essential for crafting an immersive user experience.
@@
In this work, we propose  EgoPoseFormer, a simple transformer-based model for multi-view egocentric pose estimation. As illustrated in Fig. 1, our method uses a two-stage framework [25] to overcome the joint invisibility challenge, and the body pose is predicted in a coarse-to-fine manner [39, 24, 38]. Specifically, the first stage of our model is the Pose Proposal Network (PPN), which is a simple 2-layer MLP that leverages the global information of the input multi-view feature maps to predict a joint’s coarse location. Similar to the recent heatmap-based methods [3, 32], the usage of global features allows our method to reason about the locations of all joints, including the invisible ones. Surprisingly, we found, with proper training techniques, this simple MLP can already outperform previous state-of-the-art methods in stereo inputs are available. Then, in the second stage, we employed a DETR-style [4] transformer, Pose Refinement Transformer (PRFormer), to predict 3D refinement offsets related to the first stage estimations by exploiting the multi-view stereo features and human kinematic information. Specifically, we embed each joint’s location and identity information into a Joint Query Token (JQT). Each JQT interacts with the multi-view image features and other JQTs through attention operations in each layer of PRFormer; subsequently, the refinement offsets are predicted from updated JQTs. Furthermore, we design a new Deformable Stereo Attention [48] to effectively process the fine-grained multi-view stereo features, which allows us to accurately estimate a joint’s 3D location. In summary, we make three contributions:
@@ 4.3 Ablation Studies and Discussions
In the following, we conduct ablation studies to test the proposed modules and different design choices that can affect the performance of our model.
@@
2D Heatmap Pre-training. Intuitively, 2D heatmap pre-training should tell the model what the appearance of a joint should look like, thus it can guide the deformable stereo attention to attend to relevant features, which can further help with accurately estimating the joints’ 3D locations. In Tab. 7, we report how this pre-training influences the model performance on both the stereo UnrealEgo and the monocular SceneEgo datasets. On UnrealEgo, our pre-training improves the MPJPE of the pose proposal by 3.3mm and the final prediction by 3.1mm. However, on the SceneEgo dataset, such improvement becomes more significant, with 62.2mm for the pose proposal and 29.9mm for the final prediction. The reason is that as the SceneEgo dataset does not have stereo information, the appearance features become more important in localizing a joint. This experiment validates the effectiveness of our pre-training strategy.
@@
Error Analysis. We first study the error distributions of our method. We report the error distribution of different body parts in Fig. 5 (a), encompassing both pre-refinement and post-refinement results. Our observations reveal that, for the majority of joints, our PRFormer exhibits a substantial capacity to reduce the errors in the pose proposals. Notably, except for the challenging Foot and Toe, the median error in refined poses even falls below the quartile error of the pose proposals, validating the effectiveness of our proposed PRFormer in enhancing the overall accuracy of the pose estimation. On the other hand, it is worth noting that the Foot and Toe joints, although improved in the post-refinement phase, still exhibit considerable inaccuracy, limiting the overall accuracy of the pose estimation system. Furthermore, following [12], we visualize the error distributions in the form of Cumulative Distribution Function (CDF) for different body parts. The results are shown in Fig. 5 (b), including both the pre-refinement and post-refinement CDFs. Similar to the earlier observations, the accuracy of pose estimation is substantially improved after refinement. For example, for the Hand joint, 68% of the proposals have an error less than 50mm, and this number rises to 77% for refined estimations under the same threshold. Finally, in Fig. 6, we show the qualitative examples of the most inaccurate failure cases of our method, from which we can also observe that such cases are mostly caused by joint invisibility in both views.

### v2

# EgoPoseFormer: A Simple Baseline for Stereo Egocentric 3D Human Pose Estimation

@@ Abstract
We present EgoPoseFormer, a simple yet effective transformer-based model for stereo egocentric human pose estimation. The main challenge in egocentric pose estimation is overcoming joint invisibility, which is caused by self-occlusion or a limited field of view (FOV) of head-mounted cameras. Our approach overcomes this challenge by incorporating a two-stage pose estimation paradigm: in the first stage, our model leverages the global information to estimate each joint’s coarse location, then in the second stage, it employs a DETR style transformer to refine the coarse locations by exploiting fine-grained stereo visual features. In addition, we present a Deformable Stereo Attention operation to enable our transformer to effectively process multi-view features, which enables it to accurately localize each joint in the 3D world. We evaluate our method on the stereo UnrealEgo dataset and show it significantly outperforms previous approaches while being computationally efficient: it improves MPJPE by 27.4mm (45% improvement) with only 7.9% model parameters and 13.1% FLOPs compared to the state-of-the-art. Surprisingly, with proper training settings, we find that even our first-stage pose proposal network can achieve superior performance compared to previous arts. We also show that our method can be seamlessly extended to monocular settings, which achieves state-of-the-art performance on the SceneEgo dataset, improving MPJPE by 25.5mm (21% improvement) compared to the best existing method with only 60.7% model parameters and 36.4% FLOPs. Code is available at https://github.com/ChenhongyiYang/egoposeformer.
@@ 1 Introduction
In the rapid expansion of Virtual Reality (VR) and Augmented Reality (AR) technologies [1, 2], the capability to accurately interpret and emulate human actions becomes increasingly crucial. Central to this pursuit is the egocentric pose estimation task [11, 31, 12, 37, 36, 3, 32, 35, 26, 41, 22, 45, 14], which aims to estimate the 3D body pose from a vantage point inherent to the user, predominantly from head-mounted cameras. Its precision is pivotal in a wide range of applications, such as gaming and virtual meetings, making it essential for crafting an immersive user experience for the next generation of VR/AR systems.
@@
In this work, we propose  EgoPoseFormer, a simple transformer-based model for multi-view egocentric pose estimation. As illustrated in Fig. 1, our method uses a two-stage framework [25] to overcome the joint invisibility challenge, and the body pose is predicted in a coarse-to-fine manner [39, 24, 38]. Specifically, the first stage of our model is the Pose Proposal Network (PPN), which is a simple 2-layer MLP that leverages the global information of the input multi-view feature maps to predict a joint’s coarse location. Similar to the recent heatmap-based methods [3, 32], the usage of global features allows our method to reason about the locations of all joints, including the invisible ones. Surprisingly, we found, with proper training settings, this simple MLP can already outperform previous state-of-the-art methods in stereo inputs are available. Then, in the second stage, we employed a DETR-style [4] transformer, Pose Refinement Transformer (PRFormer), to predict 3D refinement offsets related to the first stage estimations by exploiting the multi-view stereo features and human kinematic information. Specifically, we embed each joint’s location and identity information into a Joint Query Token (JQT). Each JQT interacts with the multi-view image features and other JQTs through attention operations in each layer of PRFormer; subsequently, the refinement offsets are predicted from updated JQTs. Furthermore, we design a new Deformable Stereo Attention [48] to effectively process the fine-grained multi-view stereo features, which allows us to accurately estimate a joint’s 3D location. In summary, we make three contributions:
@@ 4.3 Ablation Studies and Discussions
Error Analysis. We first study the error distributions of our method. We report the error distribution of different body parts in Fig. 5 (a), encompassing both pre-refinement and post-refinement results. Our observations reveal that, for the majority of joints, our PRFormer exhibits a substantial capacity to reduce the errors in the pose proposals. Notably, except for the challenging Foot and Toe, the median error in refined poses even falls below the quartile error of the pose proposals, validating the effectiveness of our proposed PRFormer in enhancing the overall accuracy of the pose estimation. On the other hand, it is worth noting that the Foot and Toe joints, although improved in the post-refinement phase, still exhibit considerable inaccuracy, limiting the overall accuracy of the pose estimation system. Furthermore, following [12], we visualize the error distributions in the form of Cumulative Distribution Function (CDF) for different body parts. The results are shown in Fig. 5 (b), including both the pre-refinement and post-refinement CDFs. Similar to the earlier observations, the accuracy of pose estimation is substantially improved after refinement. For example, for the Hand joint, 68% of the proposals have an error less than 50mm, and this number rises to 77% for refined estimations under the same threshold.
@@ 5 Conclusion
Appendix 0.A Extra experiments
0.A.1 2D heatmap pre-training.
Intuitively, 2D heatmap pre-training should tell the model what the appearance of a joint should look like, thus it can guide the deformable stereo attention to attend to relevant features, which can further help with accurately estimating the joints’ 3D locations. In Tab. 11, we report how this pre-training influences the model performance on both the stereo UnrealEgo and the monocular SceneEgo datasets. On UnrealEgo, our pre-training improves the MPJPE of the pose proposal by 3.3mm and the final prediction by 3.1mm. However, on the SceneEgo dataset, such improvement becomes more significant, with 62.2mm for the pose proposal and 29.9mm for the final prediction. The reason is that as the SceneEgo dataset does not have stereo information, the appearance features become more important in localizing a joint. This experiment validates the effectiveness of our pre-training strategy.
Figure 7: Qualitative comparison between EgoPoseFormer and the UnrealEgo baseline model on the UnrealEgo dataset. Ground-truths are colored in green and predictions are colored in red.
0.A.2 Qualitative comparison with the baseline
In Fig. 7, we compare the qualitative failure cases of PPN (pose proposal), PRFormer (final prediction), and the baseline ‘UnrealEgo’ model. It shows that most errors in our model’s final prediction are caused by the joint invisibility problem (mostly in lower-body). Compared with PRFormer, PPN’s estimations are more inaccurate because they are computed using the coarse global features. On the other hand, the baseline’s performance is far from satisfactory even when the joints are captured by both cameras.
0.A.3 Qualitative failure cases
In order to gain a more intuitive insight into the failure cases of our method, we present visualizations of some of the most inaccurate results in Fig. 8. A clear observation is that the majority of these failure cases are attributed to the problem of joint invisibility. For instance, in the first example in the first column, the lower body of the wearer is entirely occluded by his upper body, leading to a substantial discrepancy in the estimated locations of the lower body joints. In another instance, illustrated in the last example of the first column, nearly half of the wearer’s body extends outside the FOV of both cameras, causing severe inaccuracy for the estimated 3D pose. These examples underscore the impact of joint invisibility in egocentric 3D pose estimation. Although our method can indeed estimate the locations of invisible joints in some cases, as we explained in the main paper, such an estimation is achieved by jointly looking at the visible joints and the background scene. However, when a large part of the wearer’s body is invisible, the estimated pose is still far from accurate. Therefore, achieving accurate localization for such joints remains an important and valuable topic for future research endeavors.
0.A.4 Dependency of two stages
Figure 9: Dependency of the accuracy of PPN and PRFormer.
Here, we conduct an ablation experiment on the UnrealEgo dataset to check PRFormer’s performance when the quality of the pose proposal varies. Specifically, we use perturbed ground truth, computed by adding Gaussian noises with different scales, to serve as pose proposals, based on which we use PRFormer to compute the refined pose estimation. We plot the dependency of the two stage’s MPJPE in Fig. 9 Left. The result suggests that although PRFormer’s performance is positively related to the accuracy of the pose proposal, the refined pose estimation is always more accurate than the initial estimation, validating the effectiveness of our PRFormer.
Appendix 0.B Potential ethical concerns
Technically, one limitation of our PPN and PRFormer is that they assume headset wearers to have a full body, resulting in poor support for estimating body poses for individuals with disabilities, particularly those who have lost parts of their bodies. We acknowledge this limitation and plan to address it in future work. Another potential negative impact of our model is related to user privacy. For example, malicious agents could misuse the technology to analyze a user’s body pose without their permission. The widespread use of such technology could also potentially lead to increased surveillance and tracking of individuals, raising ethical concerns about its use in public and private spaces. Another concern could be the risk of reinforcing biases present in the training data, which could lead to inaccuracies in pose estimation for certain demographic groups. Finally, there is the potential for dependency on this technology. For example, one may first need to buy a headset before using our model, which might reduce users’ ability to perform tasks without it, affecting their autonomy and skill development.